OpenStudy (anonymous):

If sin(−2x+2y+z)=0, find the first partial derivatives ∂z∂x and ∂z∂y at the point (0, 0, 0).

OpenStudy (anonymous):

shouldn't it be -2cos(-2x+2y+z)? plug in (0,0,0) and the answer is -2?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

@ganeshie8 , that's what i thought, but webwork says the answer is wrong....any idea why?

ganeshie8 (ganeshie8):

oh wait you want to find \(\large \dfrac{\partial z}{\partial x}\)

OpenStudy (anonymous):

yeah, i thought that's what i was doing...i'm lost

ganeshie8 (ganeshie8):

sin(−2x+2y+z)=0 cos(−2x+2y+z)*(-2 + \(\dfrac{\partial z}{\partial x}\)) = 0

ganeshie8 (ganeshie8):

plugin (0,0,0) and solve \(\dfrac{\partial z}{\partial x}\)

OpenStudy (anonymous):

what happened to the "2y" in the second part? -2 is the derivative of 2x, but shouldn't 2 be the derivative of 2y and derivative of z be 1?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

y is a constant when you're `differentiating z implicitly with respect to x`

OpenStudy (anonymous):

so for the second part, it would be (2+dz/dy)? the x component is constant and differentiate the y?

ganeshie8 (ganeshie8):

yes set that equal to 0 and you get dz/dy = -2

OpenStudy (anonymous):

ok. thanks. got both parts right now.

ganeshie8 (ganeshie8):

good, you could also work it by solving z explicitly before differentiating

ganeshie8 (ganeshie8):

sin(−2x+2y+z)=0 -2x + 2y + z = arcsin0) z = arcsin(0) + 2x - 2y

OpenStudy (anonymous):

that's a good point. thanks a bunch.